OpenStudy (anonymous):

(:

OpenStudy (anonymous):

would the set up be 2(6x+y)=6xy ?

OpenStudy (e.mccormick):

Well, you actually have several pieces of info. two positive numbers <- Need to keep that in mind. twice their sum <- If you use x and y that is 2(x+y) equals their product = xy and one number is 6 times the other number. <- Now that lets you change from x and y to just x. See if 6x=y you can replace some stuff and get it in one variable.

OpenStudy (anonymous):

wait I am not really understanding the 6 times the other number

OpenStudy (anonymous):

im having difficulties writing the equation with that

OpenStudy (anonymous):

oh so you're saying that I could just use x and have one that says 6x

OpenStudy (e.mccormick):

Hmm... Yes, but it comes out oddly... I am checking it to see.

OpenStudy (anonymous):

2(6x+x)=6x x

OpenStudy (anonymous):

You're doing well. just keep going until you find one equals 14 and the other equals 7/3

OpenStudy (e.mccormick):

Yes. Those are the answers, but because of the curve... hmm...

OpenStudy (e.mccormick):

Ah, a partial substitution allows for finding it.

OpenStudy (anonymous):

wait then what was your set up if you got 7/3 and 17..?

OpenStudy (anonymous):

wasmy equation correct

OpenStudy (e.mccormick):

See, a full substitution does: 2(x+y) = xy 2(x+6x) = x 6x 2(7x) = 7x 14x = 7x And that is only true for x=0. They want positive integers and 0 is not positive so it is an extraneous answer. But if you do this: 2(x+y) = xy 2(x+6x) = xy You can solve for y. Then use y to solve for x.

OpenStudy (anonymous):

I got 7/3 and 14; not 17

OpenStudy (anonymous):

wait how did you get from x 6x to 7x

OpenStudy (anonymous):

x.6x = 6x^2

OpenStudy (e.mccormick):

oops. I was thinking +.... hmm. Odd that it still worked.

OpenStudy (e.mccormick):

I knew my math was off somewhere. Was trying to figure out where. Had a mental slip and could not see where but felt it was wrong.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I am getting 14 but i didnt get 7/3

OpenStudy (e.mccormick):

14x = 6x^2 0 = 6x^2 - 14x 0 = 3x^2 - 7x 0 = x(3x - 7) There is the 7/3 3x - 7 = 0 3x = 7 x = 7/3

OpenStudy (e.mccormick):

So much easier when I fix the math. LOL. If you look at it graphically you can see why the answers are a bit odd to find. It is not a simple line intersection: https://www.desmos.com/calculator/cioobyvoc3 But in either case for the "one number is 6 times the other number" you get the same solutions. Just their order is different if viewed as an ordered pair.

OpenStudy (anonymous):

oh ok!

OpenStudy (e.mccormick):

Sorry about the \(x^2\) mistake. Anyhow, a partial subsitution finds the y valye directly. 2(x+y) = xy 2x+2y = xy 2x+2(6x) = xy 2x+12x = xy 14x = xy 14 = y So that is one way to confirm in this case. The graph is also another nice check. Always good to do it a few ways so you can make sure of the answer.

OpenStudy (anonymous):

ah okay tots makes sense thnx!

OpenStudy (anonymous):

That's ok!. We are humans

OpenStudy (e.mccormick):

Oh, and the 14x = xy is also true when x = 0. hehe. So it does work for both times this happens on the graph

OpenStudy (e.mccormick):

And it is always the silly little mistakes that get us.